OpenStudy (sphott51):

If y=-x^2 +4x+10 is put in form y=a(x-h)^2+k , what is the value of a+h+k? The answer is 15.. Just need to show the work..

OpenStudy (loser66):

expand y = a(x-h)^2 +k, what do you have?

OpenStudy (sphott51):

Idk how to do that..

OpenStudy (anonymous):

\[y=-x^2+4x+10\] you want it to look like \[y=a(x-h)^2+k\] right? pretty obvious that \(a=-1\) then \[h=\frac{b}{2a}\]

OpenStudy (sphott51):

Ojhh

OpenStudy (anonymous):

actually \(h=-\frac{b}{2a}\)

OpenStudy (sphott51):

I made a mistake in the question lol.. the answer is actually 15..

OpenStudy (anonymous):

in your case \[-\frac{b}{2a}\] is \[-\frac{4}{2\times (-)}=2\]

OpenStudy (sphott51):

Yup, I got that far

OpenStudy (anonymous):

then to find \(k\) take \[y=-x^2+4x+10\] and replace \(x\) by \(2\)

OpenStudy (sphott51):

22?

OpenStudy (sphott51):

The answer is suppose to be 15... So idk